OpenStudy (anonymous):

Solve the equation using the quadratic formula. 2x2 = -12x - 3

OpenStudy (anonymous):

\[2x^2=-12x-3\] Move everything to the LHS (Left Hand Side) so that the quadratic equals zero. \[2x^2+12x+3=0\] Now use the quadratic formula where: \[ax^2+bx+c=0\] \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

No worries. Anytime.